Evaluate the following limits

(a) Lim y"\to" -1 Fourth root of 4y^3 minus 3

(b) lim x"\to" infinity 1+x all over 1-x

(c) lim x "\to" 2 x - 2 all over 4 - x^2

(d) lim x "\to" 0 sinx - x cos x all over x^3


1
Expert's answer
2021-07-27T17:05:08-0400

(a)


"\lim\limits_{y\to -1}(\sqrt[4]{4y^3}-3)=DNE,"

does not exist.



"\lim\limits_{y\to 1}(\sqrt[4]{4y^3}-3)=\sqrt{2}-3"


(b)


"\lim\limits_{x\to \infin}\dfrac{1+x}{1-x}=\lim\limits_{x\to \infin}\dfrac{1+x}{1-x}=\lim\limits_{x\to \infin}\dfrac{\dfrac{1}{x}+\dfrac{x}{x}}{\dfrac{1}{x}-\dfrac{x}{x}}"

"=\lim\limits_{x\to \infin}\dfrac{\dfrac{1}{x}+1}{\dfrac{1}{x}-1}=\dfrac{0+1}{0-1}=-1"

(c)


"\lim\limits_{x\to 2}\dfrac{x-2}{4-x^2}=\lim\limits_{x\to 2}\dfrac{x-2}{(2-x)(2+x)}=-\lim\limits_{x\to 2}\dfrac{1}{2+x}"

"=-\dfrac{1}{2+2}=-\dfrac{1}{4}"



(d)


"\lim\limits_{x\to 0}(\sin x-x\cos x)=0-0=0"

"\lim\limits_{x\to 0}(x^3)=0"

"\big[\dfrac{0}{0}\big]"

L'Hospital's Rule

"\lim\limits_{x\to 0}\dfrac{\sin x-x\cos x}{x^3}=\lim\limits_{x\to 0}\dfrac{(\sin x-x\cos x)'}{(x^3)'}"

"=\lim\limits_{x\to 0}\dfrac{\cos x-\cos x+x\sin x}{3x^2}=\lim\limits_{x\to 0}\dfrac{\sin x}{3x}"

"=\lim\limits_{x\to 0}\dfrac{(\sin x)'}{(3x)'}=\lim\limits_{x\to 0}\dfrac{\cos x}{3}=\dfrac{1}{3}"